OpenStudy (anonymous):

The measure of Y is twice the measure of its complement Find the measure of the complement of angle Y

OpenStudy (anonymous):

let y = x its complement = 2x x + 2x = 90 3x = 90 x = 90/3 = 30 so complement of y = 2x = 2x30 = 60